# LatentLens: Revealing Highly Interpretable Visual Tokens in LLMs

### v2

@@ 3.2 Method
Concretely, LatentLens works as follows (see also Figure˜2): Let $\mathcal{C}$ be a (large) corpus of sentences and $\mathcal{M}$ be an LLM with $L$ layers. We construct a set of reference vectors $\mathcal{R}$ by encoding every sequence $d_{j}\in\mathcal{C}$ with $\mathcal{M}$ and storing the contextualized token representations $\color[rgb]{0.05078125,0.27734375,0.6328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.05078125,0.27734375,0.6328125}\mathbf{r}_{j}^{(\ell)}$} at every position $j$ of the sequence and every layer $\ell$ of the model. To analyze the latent representation $\color[rgb]{0.75,0.2109375,0.046875}\definecolor[named]{pgfstrokecolor}{rgb}{0.75,0.2109375,0.046875}\mathbf{h}_{i}^{(\ell)}$} of a visual token at position $i$ and layer $\ell$, we compute the cosine similarity between $\color[rgb]{0.75,0.2109375,0.046875}\definecolor[named]{pgfstrokecolor}{rgb}{0.75,0.2109375,0.046875}\mathbf{h}_{i}^{(\ell)}$} and every $\color[rgb]{0.05078125,0.27734375,0.6328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.05078125,0.27734375,0.6328125}\mathbf{r}_{j}^{(\ell)}$$\in\mathcal{R}$, obtain the top-$k$ reference vectors with the highest similarity, and return their corresponding sequences as descriptions. Following the example in Figure 2, a visual token representation $\color[rgb]{0.75,0.2109375,0.046875}\definecolor[named]{pgfstrokecolor}{rgb}{0.75,0.2109375,0.046875}\mathbf{h}_{i}^{(\ell)}$} may have a nearest neighbor $\color[rgb]{0.05078125,0.27734375,0.6328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.05078125,0.27734375,0.6328125}\mathbf{r}_{j}^{(\ell)}$}, of the contextualized representation of the token clocks in the sentence: “stone tower with gold clocks”. Compared to EmbeddingLens and LogitLens, LatentLens can naturally be applied at every layer of a model and provides full sentence descriptions, allowing for more fine-grained interpretations. [^4] While we focus on sentence-level descriptions, our approach can be trivially extended to phrases or even entire paragraphs.
@@ 4.1 Experimental setup
Captioning performance. We verify that our models produce reasonable captions using DCScore (Ye et al., 2025), a GPT-4o-based judge that rates captions on a 1–10 scale across fine-grained criteria (faithfulness, detail accuracy, hallucinations, completeness). Our models average 6.0/10, with CLIP and SigLIP encoders reaching an average of 6.8 while DINOv2 models score lower at avg. 4.4, likely due to DINOv2’s lack of language supervision during pre-training. For reference, off-the-shelf Qwen2-VL-7B-Instruct (Wang et al., 2024) achieves a score of 8.5/10. See Appendix˜K.
LatentLens setup. We use 2.99M Visual Genome captions (Krishna et al., 2017) as our corpus $\mathcal{C}$ of descriptions. All captions are encoded by each individual LLM, and we store all contextual token representations for layers $\ell\in\{1,2,4,8,16,24,L\text{-}2,L\text{-}1\}$ as reference vectors $\color[rgb]{0.05078125,0.27734375,0.6328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.05078125,0.27734375,0.6328125}\mathbf{r}_{j}^{(\ell)}$}. [^5] Due to memory constraints we store at most 20 different contextual representations per token in the model’s vocabulary.
@@ 4.2 Visual token representations are consistently interpretable across layers
Results. Figure˜3 shows the fraction of interpretable visual token representations according to LatentLens, EmbeddingLens, and LogitLens for all models and layers. For EmbeddingLens (3a), we find that a reasonably large number of visual tokens are interpretable for some models, e.g., for all Olmo-based variants, 40–60% of visual tokens are interpretable from the input layer onward. For Qwen2-based models, however, only a small fraction of visual tokens are interpretable across layers (less than 20%). Llama3-based models fall in between with 20–40% of the visual tokens being labelled as interpretable.
For LogitLens (3b), we find that for all models, less than $20\%$ of visual token representations are labelled as interpretable at lower layers. For all models except Llama3 + DinoV2, Llama3 + SigLIP, and Qwen2 + SigLIP much more visual token representations are interpretable at the later layers, e.g., 60–80% for all OLMo-based models from layer 24 onward. This is consistent with the limitation of LogitLens discussed in Section˜3, in particular its applicability only for layers close to the output layer.
In contrast, using LatentLens (3c) results in consistently higher interpretability scores across all layers. Moreover, there are only minor differences in interpretability scores across model combinations: almost all models have interpretable visual tokens ranges from 60–80% across all layers (Qwen2 + CLIP is the only exception with a lower fraction of visual tokens labeled as interpretable). We perform a series of ablations (see Appendix˜D for details), confirming that our findings are not tied to our specific training setup. We find no substantial change in LatentLens interpretability when reducing the expressivity of the mapping to linear, or when training on much shorter captions.
@@ Results.
Overall, these results suggests that the visual token representations align the most with more contextualized LLM representations rather than lexical representations at the input level. We describe this as the Mid-Layer Leap phenomenon and provide additional analyses of this finding in Appendices˜F and E, investigating how much visual token representations change across layers and whether rogue dimensions (Timkey & van Schijndel, 2021) dominate the cosine similarity. We find that visual token representations change very little throughout layers, and no systematic evidence for rogue dimensions, respectively.
@@
Figure 6: Qualitative examples. Highest-scoring descriptions are extracted from different combinations of LLMs and Vision Encoders using both LatentLens and LogitLens. Each image is shown in its patchified format, where the green patches are expected to be interpretable according to the automatic judge. The described patch is shown with a red outline, and the magnitude of scores are shown in parentheses. For LatentLens, the contextualized token used for the similarity calculation is shown in bold. Best viewed in colour.
[…]
We also analyze which contextualized token representations the visual token representations are most similar (results are in Figure˜12). We obtain similar results as for the controlled setup, finding that visual token representations at layer 0 align most to contextualized text representations from layer 4, but from layer 1 onward visual token representations align mostly with contextualized text representations from the same layer (forming a diagonal pattern).
@@ 5 Qualitative results
LatentLens vs. LogitLens descriptions. Figure˜6 shows qualitative examples of the highest-scoring descriptions for two image patches extracted using LatentLens and LogitLens (see Appendix˜L and our interactive demo for additional examples). LatentLens descriptions are semantically more meaningful than those of LogitLens. For example, even when LogitLens yields some interpretable tokens at a late layer on a patch of a church tower (first row), LatentLens provides highly accurate nearest neighbors across all layers such as “large tower with clocks”. We also note that for LatentLens, the magnitude of the cosine similarity typically increases at deeper layers, i.e., the visual token representations become more similar to their textual nearest neighbors. For LogitLens, on the other hand, a higher similarity score (logit) does not necessarily translate into more interpretable descriptions. Furthermore, LogitLens descriptions often include unrenderable tokens, subwords, punctuation, and unrelated non-English tokens in Chinese. With LatentLens, however, we can simply merge the top matching token into a full word with adjacent subwords (since we have access to the entire sentence), if necessary, such as “b + elf + ry” into “belfry”. Overall, these examples highlight the advantage of LatentLens: full-word and even sentence-level descriptions are more interpretable than subword tokens. We quantify this advantage of full-sentence interpretations in Appendix˜I.
@@ 6 Related Work
Connecting (frozen) vision and language models. Several works have shown that LLMs can easily be adapted to process multimodal inputs (Tsimpoukelli et al., 2021; Mañas et al., 2023; Merullo et al., 2023; Lu et al., 2022), e.g., via small MLP or attention-based modules. Current state-of-the art VLMs (Yang et al., 2025; Deitke et al., 2025; Li et al., 2025) follow the same mapping paradigm but include more sophisticated image token pre-processing, adaptation training stages or connector designs. Some works, in spirit related to us, have also explored representing visual tokens explicitly as a weighted sum of the LLM vocabulary (Liao et al., 2025; Masry et al., 2025), instead of a vanilla MLP projector. Various works study why an LLM can so easily adapt to non-linguistic inputs: Lu et al. (2022) frame LLMs as “universal computation engines” that can process any data sequence with minimal weight updates. Patel & Pavlick (2022) argue how LLMs, only trained on language, nonetheless learn implicit world models of the physical world, e.g., of color RGB spaces. Han et al. (2025) disentangle the visual priors learned during LLM’s text-only pre-training into perception and reasoning priors. Our setup of understanding the frozen alignment between vision and language models is most similar to (Merullo et al., 2023), (Lu et al., 2022) and (Shukor & Cord, 2024), all of which explicitly keep the LLM frozen to characterize how vision integration is nonetheless possible.
@@
Interpreting VLMs. Various interpretability methods, often initially developed for unimodal text or vision models (nostalgebraist, 2020; Cunningham et al., 2023; Belinkov, 2022), have been applied to VLMs (Neo et al., 2025), e.g., to characterize cross-modal concepts (Papadimitriou et al., 2025), cross-modality circuits (Nikankin et al., 2025), and the role of attention mechanisms in extracting information from visual tokens (Neo et al., 2025; Zhang et al., 2025). Most closely related to us are works that ask what a given visual token encodes in VLMs. Aside from training-based approaches like probes on visual tokens (Fu et al., 2025), many approaches leverage the inherent embedding space of the LLM to interpret visual tokens: E.g., Mokady et al. (2021) and Jiang et al. (2025a) study whether the LLM embedding matrix can interpret visual tokens, but only show qualitative examples on 1-2 models. LogitLens has been more widely used to interpret visual tokens at later LLM layers and applied to reduce hallucinations (Neo et al., 2025; Jiang et al., 2025b; Shukor & Cord, 2024; Park & Li, 2025; Wu et al., 2025). However, these works often involve either only small studies to quantify the interpretability via LogitLens, or only few models and with a closed set of object classes as labels. Overall, neither of these works compare the interpretability via the embedding and unembedding matrix (Logitlens). To the best of our knowledge, no prior work has considered contextual embeddings for this purpose. Notably, Phukan et al. (2025) leverage the average intermediate contextual embedding of the generated answer to mitigate hallucination on VQA. They do not investigate the interpretability of visual tokens, and only rely on contextual embeddings from a single generated (not a large collection).
@@
Another fundamental question is how vision and language embedding spaces relate to each other, such as characterizing a fundamental “modality gap” (Liang et al., 2022; Jiang et al., 2024) or narrow-cone effects (Shukor & Cord, 2024) in models. Our findings can also broadly be seen as further evidence for a high structural similarity of vision and language representation spaces, coined as the Platonic Representation Hypothesis (Huh et al., 2024).
@@ 7 Discussion and Conclusion
We foresee many exciting avenues for future work. A fundamental question is to what extent vision and language representations share deeper structural similarities beyond interpretability. LatentLens may extend beyond VLMs to other non-linguistic tokens such as soft prompts (Lester et al., 2021), latent thinking (Hao et al., 2025), or speech (Ògúnrèmí et al., 2025). It could also be applied to natively multimodal models (Team, 2024; Zhou et al., 2025; Team et al., 2023) or refined as a tool with a dynamic corpus [^9] We present an initial exploration of this in Appendix J. In terms of downstream implications, we see promise in mitigating hallucination (Jiang et al., 2025b), as well as causal ablations investigating how interpretable vs. non-interpretable visual tokens affect task performance.
@@ Appendix A Limitations
  • Noun bias: The dominance of nouns in our nearest-neighbor results (approximately 45–50% as shown in Section˜H.2) may partly reflect corpus biases in Visual Genome rather than fundamental properties of visual token representations. Visual Genome’s region descriptions naturally emphasize objects and entities over actions or relations.
@@ E L2 Norm Analysis — Visual tokens have higher L2 norms and behave differently
F Layer Alignment Details — Token drift through LLM processing
G Off-the-Shelf Model Analysis — Qwen2-VL-7B-Instruct results
H Fine-grained Interpretation Analysis — POS tags, visual attributes, interpretation types
I Phrase-Level Interpretations — How helpful phrase-level interpretations are compared to single words
J Dynamic Corpus Generation — Evolutionary search to generate better descriptions with maximal cosine similarity
K Captioning Quality — Ensuring our models produce good captions with DCScore
L Qualitative Examples — 20 non-cherry-picked for LatentLens, LogitLens and EmbeddingLens
M Behind The Scenes — Beyond what the reader sees in the final product (this paper) we show stories, lessons learned, and what was discarded

### v3

@@ 3.2 Method
Concretely, LatentLens works as follows (see also Figure˜2): Let $\mathcal{C}$ be a (large) corpus of sentences and $\mathcal{M}$ be an LLM with $L$ layers. We construct a set of reference vectors $\mathcal{R}$ by encoding every sequence $d_{j}\in\mathcal{C}$ with $\mathcal{M}$ and storing the contextualized token representations $\color[rgb]{0.05078125,0.27734375,0.6328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.05078125,0.27734375,0.6328125}\mathbf{r}_{j,t}^{(\ell)}$} at every position $t$ of the sequence and every layer $\ell$ of the model. To analyze the latent representation $\color[rgb]{0.75,0.2109375,0.046875}\definecolor[named]{pgfstrokecolor}{rgb}{0.75,0.2109375,0.046875}\mathbf{h}_{i}^{(\ell^{\prime})}$} of a visual token at position $i$ and layer $\ell^{\prime}$, we compute the cosine similarity between $\color[rgb]{0.75,0.2109375,0.046875}\definecolor[named]{pgfstrokecolor}{rgb}{0.75,0.2109375,0.046875}\mathbf{h}_{i}^{(\ell^{\prime})}$} and every $\color[rgb]{0.05078125,0.27734375,0.6328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.05078125,0.27734375,0.6328125}\mathbf{r}_{j,t}^{(\ell)}$$\in\mathcal{R}$, obtain the top-$k$ reference vectors with the highest similarity, and return their corresponding sequences as descriptions. We note that, in contrast to EmbeddingLens and LogitLens, LatentLens uses several reference vectors per description, i.e., one for each token of the description and layer of the model. Therefore $\ell$ and $\ell^{\prime}$ are not necessarily equal and a top-$k$ reference vector can come from a different layer than the visual token under inspection.
Following the example in Figure 2, a visual token representation $\color[rgb]{0.75,0.2109375,0.046875}\definecolor[named]{pgfstrokecolor}{rgb}{0.75,0.2109375,0.046875}\mathbf{h}_{i}^{(\ell^{\prime})}$} may have a nearest neighbor $\color[rgb]{0.05078125,0.27734375,0.6328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.05078125,0.27734375,0.6328125}\mathbf{r}_{j,t}^{(\ell)}$}, of the contextualized representation of the token clocks in the sentence: “stone tower with gold clocks”. Compared to EmbeddingLens and LogitLens, LatentLens can naturally be applied at every layer of a model and provides full sentence descriptions, allowing for more fine-grained interpretations. [^4] While we focus on sentence-level descriptions, our approach can be trivially extended to phrases or even entire paragraphs.
@@ 4.1 Experimental setup
Captioning performance. We verify that our models produce reasonable captions using DCScore (Ye et al., 2025), a GPT-4o-based judge that rates captions on a 1–10 scale across fine-grained criteria (faithfulness, detail accuracy, hallucinations, completeness). Our models average 6.0/10, with CLIP and SigLIP encoders reaching an average of 6.8 while DINOv2 models score lower at avg. 4.4, likely due to DINOv2’s lack of language supervision during pre-training. For reference, off-the-shelf Qwen2-VL-7B-Instruct (Wang et al., 2024) achieves a score of 8.5/10. See Appendix˜L.
LatentLens setup. We use 2.99M Visual Genome captions (Krishna et al., 2017) as our corpus $\mathcal{C}$ of descriptions. All captions are encoded by each individual LLM, and we store all contextual token representations for layers $\ell\in\{1,2,4,8,16,24,L\text{-}2,L\text{-}1\}$ as reference vectors. [^5] Due to memory constraints we store at most 20 different contextual representations per token in the model’s vocabulary.
@@ 4.2 Visual token representations are consistently interpretable across layers
Results. Figure˜3 shows the fraction of interpretable visual token representations according to LatentLens, EmbeddingLens, and LogitLens for all models and layers. For EmbeddingLens (3a), we find that a reasonably large number of visual tokens are interpretable for some models, e.g., for all OLMo-based variants, 34–62% of visual tokens are interpretable from the input layer onward. For Qwen2-based models, however, only a small fraction of visual tokens are interpretable across layers (less than 20%). Llama3-based models fall in between with 20–50% of the visual tokens being labelled as interpretable.
For LogitLens (3b), we find that for all models, less than $25\%$ of visual token representations are labelled as interpretable at lower layers. For all models except Llama3 + DINOv2, Llama3 + SigLIP, and Qwen2 + SigLIP much more visual token representations are interpretable at the later layers, e.g., 60–80% for all OLMo-based models at layers 24 and 30. This is consistent with the limitation of LogitLens discussed in Section˜3, in particular its applicability only for layers close to the output layer.
In contrast, using LatentLens (3c) results in consistently higher interpretability scores across all layers. Moreover, there are only minor differences in interpretability scores across model combinations: almost all models have interpretable visual tokens in the range of 60–85% across all layers, with some OLMo variants dropping to around 53% at the final layer. We perform a series of ablations (see Appendix˜D for details), confirming that our findings are not tied to our specific training setup. We find no substantial change in LatentLens interpretability when reducing the expressivity of the mapping to linear, or when training on much shorter captions.
@@ Results.
Overall, these results suggests that the visual token representations align the most with more contextualized LLM representations rather than lexical representations at the input level. We describe this as the Mid-Layer Leap phenomenon and provide additional analyses of this finding in Appendices˜G and E, investigating how much visual token representations change across layers and whether rogue dimensions (Timkey & van Schijndel, 2021) dominate the cosine similarity. We find that visual token representations change very little throughout layers, and no systematic evidence for rogue dimensions, respectively.
@@
Figure 6: Qualitative examples. Highest-scoring descriptions are extracted from different combinations of LLMs and Vision Encoders using both LatentLens and LogitLens. The described patch is shown with a red outline, and the magnitude of scores are shown in parentheses. For LatentLens, the contextualized token used for the similarity calculation is shown in bold. Best viewed in colour.
Figure˜5 shows the fraction of interpretable visual token representations across layers for EmbeddingLens, LogitLens, and LatentLens. We again find that LatentLens leads to high interpretability (49497070%) across all layers, demonstrating that our method generalizes to an off-the-shelf model. For EmbeddingLens and LogitLens, only a small fraction of visual token representations are labeled as interpretable.
We also analyze which contextualized token representations the visual token representations are most similar (results are in Figure˜13). We obtain similar results as for the controlled setup, finding that visual token representations at layer 0 align most to contextualized text representations from layer 4, but from layer 1 onward visual token representations align mostly with contextualized text representations from the same layer (forming a diagonal pattern).
@@ 5 Qualitative results
LatentLens vs. LogitLens descriptions. Figure˜6 shows qualitative examples of the highest-scoring descriptions for two image patches extracted using LatentLens and LogitLens (see Appendix˜M and our interactive demo for additional examples). LatentLens descriptions are semantically more meaningful than those of LogitLens. For example, even when LogitLens yields some interpretable tokens at a late layer on a patch of a church tower (first row), LatentLens provides highly accurate nearest neighbors across all layers such as “large tower with clocks”. We also note that for LatentLens, the magnitude of the cosine similarity typically increases at deeper layers, i.e., the visual token representations become more similar to their textual nearest neighbors. For LogitLens, on the other hand, a higher similarity score (logit) does not necessarily translate into more interpretable descriptions. Furthermore, LogitLens descriptions often include unrenderable tokens, subwords, punctuation, and unrelated non-English tokens in Chinese. With LatentLens, however, we can simply merge the top matching token into a full word with adjacent subwords (since we have access to the entire sentence), if necessary, such as “b + elf + ry” into “belfry”. Overall, these examples highlight the advantage of LatentLens: full-word and even sentence-level descriptions are more interpretable than subword tokens. We quantify this advantage of full-sentence interpretations in Appendix˜J.
@@ 6 Related Work
Connecting (frozen) vision and language models. Several works have shown that LLMs can easily be adapted to process multimodal inputs (Tsimpoukelli et al., 2021; Mañas et al., 2023; Merullo et al., 2023; Lu et al., 2022), e.g., via small MLP or attention-based modules. Current state-of-the art VLMs (Yang et al., 2025; Deitke et al., 2025; Li et al., 2025) follow the same mapping paradigm but include more sophisticated image token pre-processing, adaptation training stages or connector designs. Some works, in spirit related to us, have also explored representing visual tokens explicitly as a weighted sum of the LLM vocabulary (Liao et al., 2025; Masry et al., 2025) or as discrete tokens (Chan et al., 2024), instead of a vanilla MLP projector. Various works study why an LLM can so easily adapt to non-linguistic inputs: Lu et al. (2022) frame LLMs as “universal computation engines” that can process any data sequence with minimal weight updates. Patel & Pavlick (2022) argue how LLMs, only trained on language, nonetheless learn implicit world models of the physical world, e.g., of color RGB spaces. Han et al. (2025) disentangle the visual priors learned during LLM’s text-only pre-training into perception and reasoning priors. Our setup of understanding the frozen alignment between vision and language models is most similar to (Merullo et al., 2023), (Lu et al., 2022) and (Shukor & Cord, 2024), all of which explicitly keep the LLM frozen to characterize how vision integration is nonetheless possible.
@@
Interpreting VLMs. Various interpretability methods, often initially developed for unimodal text or vision models (nostalgebraist, 2020; Cunningham et al., 2023; Belinkov, 2022), have been applied to VLMs (Neo et al., 2025), e.g., to characterize cross-modal concepts (Papadimitriou et al., 2025), cross-modality circuits (Nikankin et al., 2025), and the role of attention mechanisms in extracting information from visual tokens (Neo et al., 2025; Zhang et al., 2025). Most closely related to us are works that ask what a given visual token encodes in VLMs. Aside from training-based approaches like probes on visual tokens (Fu et al., 2025), many approaches leverage the inherent embedding space of the LLM to interpret visual tokens: E.g., Mokady et al. (2021) and Jiang et al. (2025a) study whether the LLM embedding matrix can interpret visual tokens, but only show qualitative examples on 1-2 models. LogitLens has been more widely used to interpret visual tokens at later LLM layers and applied to reduce hallucinations (Neo et al., 2025; Jiang et al., 2025b; Shukor & Cord, 2024; Park & Li, 2025; Wu et al., 2025). However, these works often involve either only small studies to quantify the interpretability via LogitLens, or only few models and with a closed set of object classes as labels. Overall, neither of these works compare the interpretability via the embedding and unembedding matrix (Logitlens). To the best of our knowledge, no prior work has considered contextual embeddings for this purpose. Notably, Phukan et al. (2025) leverage the average intermediate contextual embedding of the generated answer to mitigate hallucination on VQA. They do not investigate the interpretability of visual tokens, and only rely on contextual embeddings from a single generated output (not a large collection). Past work also leverages a pool of text descriptions to interpret image tokens or contributions of model components in CLIP (Chen et al., 2024; Gandelsman et al., 2024).
@@
Another fundamental question is how vision and language embedding spaces relate to each other, such as characterizing a fundamental “modality gap” (Liang et al., 2022; Jiang et al., 2024) or narrow-cone effects (Shukor & Cord, 2024) in models. Our findings can also broadly be seen as further evidence for a high structural similarity of vision and language representation spaces (Li et al., 2024), coined as the Platonic Representation Hypothesis (Huh et al., 2024).
@@ 7 Discussion and Conclusion
We foresee many exciting avenues for future work. A fundamental question is to what extent vision and language representations share deeper structural similarities beyond interpretability. LatentLens may extend beyond VLMs to other non-linguistic tokens such as soft prompts (Lester et al., 2021), latent thinking (Hao et al., 2025), or speech (Ògúnrèmí et al., 2025). It could also be applied to natively multimodal models (Team, 2024; Zhou et al., 2025; Team et al., 2023) or refined as a tool with a dynamic corpus [^9] We present an initial exploration of this in Appendix K. In terms of downstream implications, we see promise in mitigating hallucination (Jiang et al., 2025b), as well as causal ablations investigating how interpretable vs. non-interpretable visual tokens affect task performance.
@@ Appendix A Limitations
  • Noun bias: The dominance of nouns in our nearest-neighbor results (approximately 45–50% as shown in Section˜I.2) may partly reflect corpus biases in Visual Genome rather than fundamental properties of visual token representations. Visual Genome’s region descriptions naturally emphasize objects and entities over actions or relations.
@@ E L2 Norm Analysis — Visual tokens have higher L2 norms and behave differently
F Cosine Similarity Distributions — Bimodal patterns in NN similarity
G Layer Alignment Details — Token drift through LLM processing
H Off-the-Shelf Model Analysis — Qwen2-VL-7B-Instruct results
I Fine-grained Interpretation Analysis — POS tags, visual attributes, interpretation types
J Phrase-Level Interpretations — How helpful phrase-level interpretations are compared to single words
K Dynamic Corpus Generation — Evolutionary search to generate better descriptions with maximal cosine similarity
L Captioning Quality — Ensuring our models produce good captions with DCScore
M Qualitative Examples — 20 non-cherry-picked for LatentLens, LogitLens and EmbeddingLens
N Behind The Scenes — Beyond what the reader sees in the final product (this paper) we show stories, lessons learned, and what was discarded